x=10,y=8 then x+y=18 s0 x-y=?

Grade:6

9 Answers

charan
8 years ago
2
HARINATH
8 years ago
2
RAJU
8 years ago
10-8=2
SHANMUKESHWAR
8 years ago
2
SHAIK HAFEEZUL KAREEM
8 years ago
10-8=2 since they given the values of x and y
it’s very easy think    please approve my answer ….….….….….….…..thanks
G Maheshbabu
7 years ago
Here given ‘X’ value as ‘10’ and ‘Y’ value as ‘8’ and X+Y=10.then “X-Y=2” is the answer.It can be check in in this way (X+Y)*(X-Y)=(X^2-y^2),Substitute X+Y and X-Y values, we get answer as “2”
kalyan
7 years ago
answer is 2 because x=10 and y=8 so x-y= 2,, and x+y = 18 and also x*y=80 this is answer of this question    and also i know one formula to find X-y value to use x^2 and y^2 valuea and x+y value … formula is (x-y)=(x^2*y^2)/(x+y)
kalyan
7 years ago
answer is 2 because x=10 and y=8 so x-y= 2,, and x+y = 18 and also x*y=80 this is answer of this question
SREEKANTH
7 years ago
method 1:-the value of x=10
the value of y=8
  x-y=10-8=2
so the answer for this question is 2
method:-x-y=(x+y)^2-x^2-y^2-2xy+x-y=324-64-100-160=10-8=2 so the answer for this question is 2